A company produces light bulbs whose life follows a normal distribution with a mean of 1,200

hours and standard deviation 200 hours. If we choose a light bulb at random, what is the

probability that its lifetime will be between 1100 and 1,600 hours?


1
Expert's answer
2021-06-03T15:33:14-0400

"\mu=1200"

"\sigma=200"



"P(1100<X<1600)="

"=P(\frac{1100-\mu}{\sigma}<Z<\frac{1600-\mu}{\sigma})="

"=P(\frac{1100-1200}{200}<Z<\frac{1600-1200}{200})=P(-0.5<Z<2)="

"=P(Z<2)-P(Z<-0.5)=0.9772-0.3085=0.6687"


Answer: 0.6687 is probability that its lifetime will be between 1100 and 1,600 hours